[Hallc_running] [New Logentry] RC Daily Update

bduran at jlab.org bduran at jlab.org
Sat Feb 15 17:45:01 EST 2020

Logentry Text:
Yesterday, during the opportunistic controlled access  power supply for holding field is replaced with the new Kepco. Following that NMR stopped working (https://logbooks.jlab.org/entry/3783977) and unfortunately we lost all the target polarization. We had another controlled access later to replace the power supply with the old one.  We took carbon optics (sieve in) data while waiting for the target to reach about 40% polarization to switch to the production running. While we had beam back we were having position lock problems (https://logbooks.jlab.org/entry/3784013) which seemed to be okay soon after (didn't last too long though).
Since late swing shift the beam quality have been very poor due to several BLM trips happening after bunch compression (https://logbooks.jlab.org/entry/3784312) and the beam positions mostly looked unstable. It's hard to say if the positions look awkward while they're ramping up the current or not as the beam was tripping very frequently. At today's MCC meeting I made it clear that this is concerning for us, especially when we start running with polarized 3He target. It seems the change they made afterwards on 0L04 phase worked as the positions were pretty stable since then (https://logbooks.jlab.org/entry/3784402 https://logbooks.jlab.org/entry/3784312).

Around 8:30 am in the morning NMR measurement gave 41% longitudinal polarization and we started to set up for the production running. We switched to transverse and did an NMR before data taking which gave 36.9% transverse polarization (https://logbooks.jlab.org/entry/3784363).We took 6 hrs of 3He transverse data and did an NMR measurement in the middle: 46.4% transverse polarization (https://logbooks.jlab.org/entry/3784451). At 4 pm we switched to longitudinal and made an NMR (53% longitudinal) (https://logbooks.jlab.org/entry/3784538) We have been taking data with this spin configuration since then.

The beam will be gone today around 7 pm for a couple of hours due to Hall B pass change. We will take an opportunistic controlled access at that time to do laser work (plate adjustment) in the hall. Once we get the beam back probably during owl shift we will continue with production data taking.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3784578
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200215/7dc63944/attachment.html>

More information about the Hallc_running mailing list